# bibliography.bib

### v1

@@ 3.1 Notable examples of Raney extensions
For a frame $L$, there is an adjunction
[…]
$F\mapsto\bigcap\varphi[F],$
[…]
@@
[…]
@@
For any spatial frame $L$, the pair $(L,\mathcal{U}(\mathsf{pt}(L)))$ is a Raney extension, where we have identified each element $a\in L$ with the open $\varphi(a)$.
@@
For a spatial frame $L$, the pair $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op})$ is a Raney extension. By Theorem 3.8, there is an isomorphism $\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op}\cong\mathcal{U}(\mathsf{pt}(L))$, and by definition of the isomorphism this maps principal filters to opens of $\mathsf{pt}(L)$. ∎
Let us now consider the motivating example behind the definition of Raney extension.
@@
For every topological space $X$ the Raney extension $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{CP}$-dense.
@@
Let $X$ be a topological space. Consider some family $U_{i}\subseteq X$ of open sets. To show the claim, we use Corollary 3.3, and show that ${\uparrow}^{\Omega(X)}\bigcap_{i}U_{i}$ is an intersection of completely prime filters in $\mathcal{U}(X)^{*}$. The filter may be written as
$\bigcap\{N(x):x\in\bigcap_{i}U_{i}\},$
where $N(x)$ denotes the neighborhood filter of a point $x\in X$. Indeed, a neighborhood filter $N(x)$ is $\{U\in\Omega(X):{\uparrow}x\subseteq U\}$, and so it is in $\mathcal{U}(X)^{*}$. ∎
@@
Suppose that $X$ is a sober space, and let $P$ be a completely prime filter. By definition of $\mathcal{CP}$-compactness, we have to show that $\bigcap P\subseteq U$ implies that $U\in P$ for every open $U\subseteq X$. By sobriety, each completely prime filter is the neighborhood filter of some point, and so we may assume $P=N(x)$ for some $x\in X$. We have $\bigcap N(x)={\uparrow}x$, and so, indeed, $\bigcap N(x)\subseteq U$ implies that $x\in U$, and so $U\in N(x)$. For the second part of the claim, assume that for some $T_{0}$ space $X$ the pair $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{CP}$-compact. Let us show that in $X$ all completely prime filters are neighborhood filters. Let $P\subseteq\Omega(X)$ be a completely prime filter. By the characterization in Proposition 3.2, this is of the form $\{U\in\Omega(X):Y\subseteq U\}$, for some saturated set $Y$. By complete primality, the set $Y$ is then completely join-prime in the lattice $Y$, and so it is a principal upset ${\uparrow}x$ for some $x\in X$. Hence $P=N(x)$. This means that $X$ is sober. The last part of the claim follows by combining this characterization of sobriety with Lemma 3.10. ∎
4 Extensions of frame maps
We ask when a map $f:L\to M$ of frames can be extended to Raney extensions of these frames. For a collection of filters $\mathcal{F}$ and an arbitrary filter $G$ of a frame $L$, we introduce the following closure operator in $\mathsf{Filt}(L)$:
$\mbox{cl}_{\mathcal{F}}(G)=\bigcap\{F\in\mathcal{F}:G\subseteq F\}.$
In case $\mathcal{F}=\mathsf{Filt}_{\mathcal{SO}}(L)$, as an abbreviation we will denote this closure operator as $\mbox{cl}_{\mathcal{SO}}$, and similarly for similarly named collections of filters. We note that for a subcolocale $\mathcal{F}\subseteq\mathsf{Filt}(L)^{op}$, the meet of a family $F_{i}$ is computed as $\mbox{cl}_{\mathcal{F}}(\bigcup_{i}F_{i})$. Recall that, by Theorem 3.1, we may identify elements of the coframe components of Raney extensions with filters of the form ${\uparrow}^{L}a$. Then, if we are seeking to extend a frame map $f:L\to M$ to Raney extensions $(L,C)$ and $(M,D)$, we are looking for a coframe morphism making the following commute:
[…]
By commutativity of the diagram, if this map exists, it has to act as ${\uparrow}a\mapsto{\uparrow}f(a)$ for each $a\in L$. Because it also has to be meet-preserving, and because the order on $C^{*}$ and $D^{*}$ is reverse set inclusion, for a filter $F\in C^{*}$ it has to act as $\bigcup\{{\uparrow}f:f\in F\}\mapsto\mbox{cl}_{D^{*}}(\bigcup\{{\uparrow}f(x):x\in F\})$. This means that if such lifting exists, then it has to be defined as $F\mapsto\mbox{cl}_{D^{*}}(f[F])$.
@@ 4 Extensions of frame maps
Suppose that we have coframes $C$ and $D$, such that $L\subseteq C$ is a subframe that meet-generates $C$. Suppose that a map $f:C\rightarrow D$ preserves all meets, as well as all the joins of $L$. Then it is a coframe map.
@@
To show that it is a coframe map, it suffices to show that it preserves finite joins. Consider $x,y\in C$. We have $x=\bigwedge_{i}x_{i}$ and $y=\bigwedge_{j}y_{j}$ for families $x_{i},y_{j}\in L$. Then
$f(x\vee y)=f(\bigwedge_{i}x_{i}\vee\bigwedge_{j}y_{j})=\bigwedge_{i,j}f(x_{i})\vee f(y_{j})=f(x)\vee f(y).\qed$

### v2

@@ 4 Topological duality for Raney extensions
For a frame $L$ and for a sublocale $\mathcal{F}\subseteq\mathsf{Filt}(L)$ such that it contains all principal filters, $\mathsf{pt}_{R}(\mathcal{F}^{op})=\mathsf{Filt}_{\mathcal{CP}}(L)\cap\mathcal{F}$.
@@
We show that an element $P\in\mathcal{F}$ is completely prime in the frame $\mathcal{F}$ if and only if it is completely prime as an element of $\mathsf{Filt}(L)$. If an element $P\in\mathcal{F}$ is completely prime in the frame $\mathsf{Filt}(L)$, then it is also completely prime as an element of $\mathcal{F}$, as meets of elements of $\mathcal{F}$ are a subset of all the meets in $\mathsf{Filt}(L)$. For the converse, suppose that $P$ is completely prime in $\mathcal{F}$, and that $\bigvee_{i}x_{i}\in P$ for some collection $x_{i}\in L$. This means $\bigcap_{i}{\uparrow}x_{i}\subseteq P$, and because $\mathcal{F}$ contains all principal filters and by assumption on $P$, ${\uparrow}x_{i}\subseteq P$ for some $i\in I$. ∎
@@
A Raney extension $(L,C^{*})$ has as points the elements of $C^{*}\cap\mathsf{Filt}_{\mathcal{CP}}(L)$, and as opens the sets of the form $\{P\in\mathsf{Filt}_{\mathcal{CP}}(L)\cap C^{*}:a\in P\}$ for some $a\in L$.
@@
The first part of the statement is a direct consequence of Theorem 4.4. For the second part of the statement, it suffices to unravel the definition of the topology on $\mathsf{pt}_{R}(L,C^{*})$. ∎
We now define the left adjoint to $\mathsf{pt}_{R}$. For a topological space $X$ we define $\Omega_{R}(X)$ as the pair $(\Omega(X),\mathcal{U}(X))$, we extend the assignment to morphisms as $f\mapsto f^{-1}$.
@@
For every Raney extension $(L,C)$ there is a surjective map of Raney extensions $\varphi_{(L,C)}:(L,C)\rightarrow\Omega_{R}(\mathsf{pt}_{R}(L,C))$. This is an isomorphism precisely when $C$ is join-generated by its completely join-prime elements.
@@
The fact that it is a surjection and a map of Raney extensions follows from properties 1 and 2 of the topologizing map $\varphi_{(L,C)}$. The map is an isomorphism precisely when it is injective, and this happens exactly when for $c,d\in C$ such that $c\nleq d$ there is some $x\in\mathsf{pt}_{R}(C)$ such that $x\leq c$ and $x\nleq d$. This holds if and only if the completely join-prime elements join-generate $C$. ∎
The map we have just defined will be the evaluation at an object of the natural transformation $\Omega_{R}\circ\mathsf{pt}_{R}\Rightarrow 1_{\mathbf{Raney}^{op}}$. Let us now define the other natural transformation $1_{\mathbf{Top}}\Rightarrow\mathsf{pt}_{R}\circ\Omega_{R}$.
Lemma 4.7.
For every topological space $X$ the map $\psi_{X}:X\rightarrow\mathsf{pt}_{R}(\Omega_{R}(X))$ defined as $x\mapsto{\uparrow}x$ is a continuous map. This is a homeomorphism precisely when $X$ is a $T_{0}$ space.
Proof.
That the map is well-defined and surjective follows from the observation that the completely join-prime elements of $\mathcal{U}(X)$ are precisely the principal upsets. For continuity, we observe that the $\psi_{X}$-preimage of an open set $\varphi(U)$ is the set $\{x\in X:{\uparrow}x\in\varphi(U)\}=U$. This map is also open, as the direct image of an open $U\subseteq X$ is the open $\{{\uparrow}x:{\uparrow}x\subseteq U\}=\varphi(U)$. The map is then a homeomorphism when it is injective, and this holds if and only if, whenever $x\neq y$, ${\uparrow}x\neq{\uparrow}y$. This amounts to the specialization preorder being an order, that is, the space being $T_{0}$. ∎
Recall that an adjunction $L:\mathcal{C}\leftrightarrows\mathcal{D}:R$ is said to be idempotent if every element of the form $R(d)$ for some object $d\in\mathsf{Obj}(\mathcal{D})$ is a fixpoint on the $\mathcal{C}$ side, and the same holds for the $\mathcal{D}$ side.
Theorem 4.8.
The pair $(\Omega_{R},\mathsf{pt}_{R})$ constitutes an idempotent adjunction $\mathbf{Top}\leftrightarrows\mathbf{Raney}^{op}$. Raney duality is the restriction of this adjunction to a dual equivalence.
Proof.
The proof of adjointness amounts to standard computations. The two maps in Lemmas 4.6 and 4.7 are the counit and the unit, respectively. Let us see that the adjunction is idempotent. By Lemma 4.6, any Raney extension $(\Omega(X),\mathcal{U}(X))$ is a fixpoint, as the coframe $\mathcal{U}(X)$ is join-generated by the elements of the form ${\uparrow}x$. By Lemma 4.7, any $T_{0}$ space is a fixpoint. ∎
Motivated by the result above and by Lemma 4.6, we say that a Raney extension $(L,C)$ is spatial if $C$ is join-generated by the completely join-prime elements.
Proposition 4.9.
A Raney extension $(L,C)$ is spatial if and only if $C^{*}\subseteq\mathcal{I}(C^{*}\cap\mathsf{Filt}_{\mathcal{CP}}(L))$.
Proof.
Because of the isomorphism ${\uparrow}^{L}:C\cong C^{*}$, a Raney extension $(L,C)$ is spatial precisely when all elements of $C^{*}$ are intersections of completely join-prime elements in $C^{*}$, as by Corollary 4.5, $\mathsf{pt}_{R}(C^{*})=\mathsf{Filt}_{\mathcal{CP}}(L)\cap C^{*}$. ∎
Corollary 4.10.
Spatial Raney extensions are $\mathcal{CP}$-dense.
Proof.
This follows from Corollary 3.3, and Proposition 4.9. ∎
4.1 The collection of Raney spectra on a frame
In this subsection, our final goal is proving that, on a frame $L$, for any Raney extension $(L,C)$ there are subspace inclusions $\mathsf{pt}_{D}(L)\subseteq\mathsf{pt}_{R}(L,C)\subseteq\mathsf{pt}(L)$.
Lemma 4.11.
For a frame $L$, for any $a\in L$ the meet $\bigwedge\{x\in L:a<x\}$ is exact.
Proof.
Let $L$ be a frame and let $a\in L$. Let us consider the meet $\bigwedge\{x\in L:a<x\}$. Let $b\in L$. We claim that $\bigwedge\{x\vee b:a<x\}\leq\bigwedge\{x\in L:a<x\}\vee b$. We consider two cases. First, let us assume that $b\leq a$. If this is the case, then $b\leq x$ whenever $a<x$, and so both the left hand side and the right hand side equal $\bigwedge\{x\in L:a<x\}$. Now, let us assume instead that $b\nleq a$. This is equivalent to saying that $a<a\vee b$. This means that:
$\bigwedge\{x\vee b:a<x\}\leq a\vee b\leq\bigwedge\{x\in L:a<x\}\vee b.\qed$
Proposition 4.12.
A completely prime filter $L{\setminus}{\downarrow}p$ is exact if and only if the prime $p$ is covered.
Suppose that the completely prime filter $L{\setminus}{\downarrow}p$ is exact. To show that the prime $p$ is covered, we prove that $\bigwedge\{x\in L:p<x\}\nleq p$. By Lemma 4.11, the meet on the left-hand side is exact. The result follows by our assumption that $L{\setminus}{\downarrow}p$ is closed under exact meets. For the converse, we suppose that $p$ is a covered prime and that $x_{i}\nleq p$ for the members of some family $\{x_{i}:i\in I\}$ such that their meet is exact. We then have that $x_{i}\vee p\neq p$ for every $i\in I$, and as $p$ is covered, this implies that $\bigwedge_{i}(x_{i}\vee p)\neq p$. By exactness of the meet $\bigwedge_{i}x_{i}$, we also have $(\bigwedge_{i}x_{i})\vee p\neq p$, that is $\bigwedge_{i}x_{i}\nleq p$, as required.
Lemma 4.13.
For any frame $L$, the spectrum of $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ is homeomorphic to the space $\mathsf{pt}_{D}(L)$. The spectrum of $(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$ is the classical spectrum $\mathsf{pt}(L)$.
Proof.
By Corollary 4.5, the points of $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ are the completely prime filters which are also exact. By Proposition 4.12, these are the filters of the form $L{\setminus}{\downarrow}p$ for some covered prime $p\in L$. Indeed, then, there is a bijection between the points of $\mathsf{pt}_{R}(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ and those of $\mathsf{pt}_{D}(L)$. This is a restriction and co-restriction of the standard homeomorphism between the spectrum $\mathsf{pt}(L)$ and its space of completely prime filters, and so it is a homeomorphism. For $(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$, it suffices to notice that since all completely prime filters are strongly exact, $\mathsf{Filt}_{\mathcal{CP}}(L)\cap\mathsf{Filt}_{\mathcal{SE}}(L)=\mathsf{Filt}_{\mathcal{CP}}(L)$. ∎
We shall now refine the result above to the case of subfit frames. We call $\mathsf{maxpt}(L)$ the collection of maximal primes of a frame $L$, equipped with the subspace topology inherited from $\mathsf{pt}(L)$.
Proposition 4.14.
Let $L$ be a frame. A prime $p\in L$ is maximal if and only if $L{\setminus}{\downarrow}p$ is a regular filter.
Proof.
Suppose that $p\in L$ is a maximal prime. Because it is maximal, ${\uparrow}p=\{p,1\}$. We claim that the completely prime filter $L{\setminus}{\downarrow}p$ is its pseudocomplement in the frame of filters. Indeed, $L{\setminus}{\downarrow}p\cap\{1,p\}=\{1\}$. Furthermore, if, for a filter $F$, $F\cap\{1,p\}=\{1\}$ then $p\notin F$, and so for $f\in F$ we must have $f\nleq p$. For the converse, suppose that $p\in L$ is such that $L{\setminus}{\downarrow}p$ is a regular filter. By Proposition 2.12, this is the intersection of a collection of filters of the form $\{x\in L:x\vee a=1\}$ for some $a\in L$. As $L{\setminus}{\downarrow}p$ is completely prime, it must be $\{x\in L:x\vee a=1\}$ for some $a\in L$. This means that for all $x\in L$ the conditions $x\leq p$ and $x\vee a\neq 1$ are equivalent. In particular, because the filter is not all of $L$ (as it is completely prime), we must have $a\leq p$ since $a\vee a=a\neq 1$. This means that if $x\nleq p$ then $x\vee a=1$ and so $x\vee p=1$, for all $x\in L$. This means that $p$ must be maximal. ∎
Proposition 4.15.
For a subfit frame $L$, the spectrum of the Raney extension $(L,\mathsf{Filt}_{\mathcal{R}}(L)^{op})$ is the $T_{1}$ space $\mathsf{maxpt}(L)$.
Proof.
Suppose that $L$ is a subfit frame. We claim that all its exact filters are regular. By Proposition 2.14, $\mathsf{Filt}_{\mathcal{R}}(L)$ contains all principal filters, and so by Lemma 3.7 we must have $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathsf{Filt}_{\mathcal{R}}(L)$. The reverse inclusion holds for all frames. By Corollary 4.5, then, the points of $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ are the regular completely prime filters, which by Proposition 4.14 are those corresponding to maximal primes of $L$. The fact that this is a homeomorphism comes from the fact that this is a restriction of the standard homeomorphism between the spectrum $\mathsf{pt}(L)$ and the spectrum defined in terms of prime elements of $L$. The space $\mathsf{maxpt}(L)$ is a $T_{1}$ space, since whenever $p,q\in\mathsf{maxpt}(L)$, both $p\nleq q$ and $q\nleq p$ by maximality, and so the open set $\{a\in L:a\nleq p\}$ contains $q$ and omits $p$, and the open set $\{a\in L:a\nleq q\}$ contains $p$ and omits $q$. ∎
Lemma 4.16.
For a Raney extension $(L,C)$ there are subspace embeddings
$\mathsf{pt}_{D}(L)\to\mathsf{pt}_{R}(L,C)\to\mathsf{pt}(L).$
Proof.
If $(L,C)$ is a Raney extension, $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}\subseteq\mathsf{Filt}_{\mathcal{SE}}(L)$, by Theorem 3.6 and by Proposition 3.8. Therefore,
$\mathsf{Filt}_{\mathcal{CP}}(L)\cap\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathsf{Filt}_{\mathcal{CP}}(L)\cap C^{*}\subseteq\mathsf{Filt}_{\mathcal{CP}}(L)\cap\mathsf{Filt}_{\mathcal{SE}}(L).$
By Corollary 4.5, this means that thre is a chain of subspace inclusions $\mathsf{pt}_{R}(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})\subseteq\mathsf{pt}_{R}(L,C)\subseteq\mathsf{pt}_{R}(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$. The result follows from Lemma 4.13. ∎
Lemma 4.17.
For a frame $L$ and a subset $\mathcal{X}\subseteq\mathsf{Filt}(L)$ the smallest sublocale $\mathcal{S}(\mathcal{X})$ is the set ($\mathcal{I}(\{{\uparrow}a\rightarrow F:a\in L,F\in\mathcal{X}\})$.
Proof.
By Lemma 2.2, it suffices to show that the collection in the claim is the same as ($\mathcal{I}(\{G\rightarrow F:G\in\mathsf{Filt}(L),F\in\mathcal{X}\})$. Indeed, for each $G\in\mathsf{Filt}(L)$ and $F\in\mathcal{X}$, $G\rightarrow F=\bigcap\{{\uparrow}g\rightarrow F:g\in G\}$. ∎
The following fact follows directly from Lemma 2.5, and the fact that completely prime filters are prime elements of $\mathsf{Filt}(L)$.
Lemma 4.18.
For a frame $L$ and a completely prime filter $P\subseteq L$, for each $a\in L$,
${\uparrow}a\rightarrow P=\begin{cases}L&\text{ if $a\in P$}\\ P&\text{ otherwise.}\end{cases}$
Theorem 4.19.
The spectra of Raney extensions over $L$ coincide, up to homeomorphisms, with the interval
$[\mathsf{pt}_{D}(L),\mathsf{pt}(L)]$
of the powerset of $\mathsf{pt}(L)$.
Proof.
The spectrum of a Raney extension $(L,C)$ is contained in the $[\mathsf{pt}_{D}(L),\mathsf{pt}(L)]$ interval by Lemma 4.16. To show the converse, because of Lemma 4.13, it suffices to show that for any collection of completely prime filters $\mathcal{P}$ such that $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathcal{P}$ there is some Raney extension $(L,C)$ such that its points are $\mathcal{P}$, that is, $C^{*}\cap\mathsf{Filt}_{\mathcal{CP}}(L)=\mathcal{P}$. Let then $\mathcal{P}$ be such a collection. Consider the sublocale $\mathcal{S}(\mathcal{P}\cup L)\subseteq\mathsf{Filt}(L)$. By Lemma 3.7, this is the same as $\mathcal{S}(\mathcal{P}\cup\mathsf{Filt}_{\mathcal{E}}(L))$. Observe that $\mathsf{Filt}_{\mathcal{E}}(L)$ is stable under ${\uparrow}a\to-$ for each $a\in L$, as it is a sublocale. The same holds for $\mathcal{P}$, by Lemma 4.18. By Lemma 4.17, $\mathcal{S}(\mathcal{P}\cup L)=\mathcal{I}(\mathcal{P}\cup\mathsf{Filt}_{\mathcal{E}}(L))$. We now consider the Raney extension $(L,\mathcal{S}(\mathcal{P}\cup L)^{op})$. It is clear that all the elements of $\mathcal{P}$ are points of this Raney extension, by Corollary 4.5. Let us show the reverse set inclusion. Suppose that there is a completely prime filter $F$ such that $F\in\mathcal{S}(\mathcal{P}\cup L)$. By complete primality, and by the characterization above, this is either in $\mathcal{P}$ or in $\mathsf{Filt}_{\mathcal{E}}(L)$. In the second case, it is in $\mathcal{P}$, too, by assumption on $\mathcal{P}$. Indeed, then, $\mathcal{S}(\mathcal{P}\cup L)\cap\mathsf{Filt}_{\mathcal{CP}}(L)=\mathcal{P}$, as desired. ∎
5 Topological properties and Raney extensions
5.1 Sobriety
We now look at sobriety and characterize it in terms of Raney extensions.
Lemma 5.1 ([jakl24], Lemma 5.4).
For every frame $L$ there is an isomorphism $\iota:\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op}\cong\mathcal{U}(\mathsf{pt}(L))$ defined on generators as $\iota(P)={\uparrow}P=\{Q\in\mathsf{Filt}_{\mathcal{CP}}(L):P\subseteq Q\}$.
Lemma 5.2.
Let $X$ be a $T_{0}$ space. For a saturated set $Y\subseteq X$, the filter ${\uparrow}^{\Omega(X)}Y$ is completely prime if and only if $Y={\uparrow}x$ for some $x\in X$.
Proof.
For all $y\in Y$, as $Y$ is saturated, ${\uparrow}y\subseteq Y$. Towards contradiction, suppose that $Y\nsubseteq{\uparrow}y$ for each $y\in Y$. For each $y\in Y$, let $y^{\prime}\in Y{\setminus}{\uparrow}y$, so $y\nleq y^{\prime}$, and let $U_{y}\in\Omega(X)$ be such that $y\in U_{y}$ and $y^{\prime}\notin U_{y}$. Then, $Y\subseteq\bigcup\{U_{y}:y\in Y\}$. But $Y\nsubseteq U_{y}$ for all $y\in Y$ as $y^{\prime}\in Y{\setminus}U_{y}$. This contradicts complete primality of ${\uparrow}^{\Omega(X)}Y$. ∎
@@ 5.1 Sobriety
Suppose that $X$ is a sober space, and let $P$ be a completely prime filter. By sobriety, we may assume $P=N(x)$ for some $x\in X$. Since $\bigcap N(x)={\uparrow}x$, indeed, $\bigcap N(x)\subseteq U$ implies that $x\in U$, and so $U\in N(x)$. Then, $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{CP}$-compact. For the second part of the claim, assume that for some $T_{0}$ space $X$ the pair $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{CP}$-compact. We show sobriety by showing that any completely prime filter $P$ is a neighborhood filter. By the characterization in Proposition 3.2, this is of the form ${\uparrow}^{\Omega(X)}Y$, for some saturated set $Y$. By Lemma 5.2, $Y={\uparrow}x$ for some $x\in X$. Hence $P=N(x)$. The last part of the claim follows by combining this characterization of sobriety with Corollary 4.10. ∎
Motivated by this result, we define a Raney extension $(L,C)$ to be sober if it is $\mathcal{CP}$-compact. In Section 6 we will show the existence of sober coreflections in $\mathbf{Raney}$. Let us now compare sobriety with spatiality for Raney extensions.
@@
A Raney extension $(L,C)$ is sober and spatial if and only if it is $\mathcal{CP}$-canonical.
@@
It follows from Proposition 4.9 and by Proposition 3.2 that a Raney extension $(L,C)$ is sober and spatial if and only if $C^{*}=\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))$. This holds if and only if the Raney extension is $\mathcal{CP}$-canonical. ∎
Proposition 5.5.
For a spatial frame $L$, $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op})$ is the unique (up to isomorphism) sober and spatial Raney extension.
Proof.
By Lemma 5.4, when a sober and spatial Raney extension exists, it is unique, up to isomorphism, by Theorem 3.6. If $L$ is a spatial frame, then $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op})$ is a Raney extension by Proposition 2.14, and it is the $\mathcal{CP}$-canonical Raney extension by Theorem 3.6. ∎
5.2 The TD axiom
Let us now look at the Raney analogue of the $T_{D}$ axiom.